### v2

# Revisiting Graph Neural Networks for Graph-level Tasks: Taxonomy, Empirical Study, and Future Directions

Abstract.
Graphs are fundamental data structures for modeling complex interactions in domains such as social networks, molecular structures, and biological systems. Graph-level tasks, which involve predicting properties or labels for entire graphs, are crucial for applications like molecular property prediction and subgraph counting. While Graph Neural Networks (GNNs) have shown significant promise for these tasks, their evaluations are often limited by narrow datasets, task coverage, and inconsistent experimental setups, hindering their generalizability. In this paper, we present a comprehensive experimental study of GNNs on graph-level tasks, systematically categorizing them into five types: node-based, hierarchical pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based GNNs. To address these challenges, we propose a unified evaluation framework OpenGLT for graph-level GNNs. OpenGLT standardizes the evaluation process across diverse datasets, multiple graph tasks (e.g., classification and regression), and real-world scenarios, including noisy, imbalanced, and few-shot graphs. Extensive experiments are conducted on 16 baseline models across five categories, evaluated on 13 graph classification and 13 graph regression datasets. These experiments provide comprehensive insights into the strengths and weaknesses of existing GNN architectures.
Graph Neural Networks, Graph-level tasks, Unified evaluation
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 9–13, 2026; Jeju, Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining (KDD ’26), August 9–13, 2026, Jeju, Koreaisbn: 978-1-4503-XXXX-X/26/08copyright: none
1. Introduction
Graphs represent objects as nodes and their relationships as edges, which are key data structures across various domains to model complex interactions (Bonifati et al., 2024; Fan, 2022; Fang et al., 2022), such as social networks, molecular structures, biological systems, etc. Early graph representation learning methods (Perozzi and others, 2014; Grover and Leskovec, 2016; Wang et al., 2016; Dong et al., 2017) laid the foundation of graph learning. Graph-level tasks aim at predicting properties of an entire graph, rather than individual nodes or edges. These tasks are crucial in domains including subgraph counting in database management (Fichtenberger and Peng, 2022; Li and Yu, 2024), molecular property prediction in chemistry (Zhong and Mottin, 2023; Hu et al., 2020a), and protein classification in bioinformatics (Morris et al., 2020; Li et al., 2024a). Recently, graph neural networks (GNNs) (Wang et al., 2023b; Xiang et al., 2025; Liu et al., 2025; Song et al., 2023; Liao et al., 2022; Cui et al., 2021; Li et al., 2024e) have emerged as powerful tools for graph-structured and anomaly detection tasks (Ma et al., 2023). They learn node representations by iteratively aggregating neighbor information to obtain graph representations.
Depending on the approach, we categorize GNNs for graph-level tasks into five categories: node-based, hierarchical pooling (HP)-based, subgraph-based, graph learning (GL)-based, and self-supervised learning (SSL)-based methods. Node-based GNNs (Kipf and Welling, 2016; Hamilton et al., 2017; Xu et al., 2019; Veličković et al., 2017) compute node representations through message passing and aggregate them via a permutation-invariant readout function, such as averaging, to form the final graph embedding. HP-based GNNs (Mesquita et al., 2020; Dhillon et al., 2007; Li et al., 2022b; Bianchi et al., 2020) apply pooling operations to reduce the graph size and capture hierarchical structure, yielding multi-level graph representations. Subgraph-based GNNs (Ding et al., 2024; Yan et al., 2024; Bevilacqua et al., 2024; Papp and Wattenhofer, 2022) divide the graph into subgraphs, learn a representation for each, and then aggregate these to represent the whole graph. GL-based GNNs (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d, b) enhance graph quality by reconstructing structure and features. SSL-based GNNs (Sun et al., 2024; Wang et al., 2022b; Inae et al., 2023; Li et al., 2024a; Thakoor et al., 2021) pretrain on unlabeled data, either by predicting graph properties or maximizing agreement between augmented views of the same graph.
Although various GNNs have been designed for graph-level tasks, their evaluations are often restricted to a narrow range of domain-specific datasets and insufficient baseline comparisons (Zhiyao et al., 2024; Errica et al., 2020; Li et al., 2023; Wang et al., 2024b). To ensure fair comparison, we identify and address three key shortcomings in current evaluation frameworks.
  • Issue 1. No clear taxonomy for GNNs on graph-level Tasks. Graph-level tasks require different approaches than node-level tasks, yet a clear taxonomy for GNNs on graph-level tasks is lacking. This gap hinders holistic understanding and systematic comparison of models.
  • Issue 2. Restricted Coverage of GNN Architectures. Most evaluations focus on a limited set of architectures, such as node-based GNNs, while often ignoring more expressive models like subgraph-based GNNs. This narrow coverage limits performance comparisons and overlooks the strengths of diverse approaches.
  • Issue 3. Insufficient Data Diversity. Current evaluations typically use datasets from a narrow range of domains, such as chemistry or biology. This limited diversity can lead to overfitting and restricts the generalizability of GNNs to other domains like social networks or different types of graphs.
  • Issue 4. Narrow Task and Scenario Scope. Current evaluation frameworks typically focus on a single type of graph-level task, such as molecular graph classification, and overlook diverse applications like cycle or path counting. They also assume access to ample clean labeled data, neglecting real-world challenges such as noise, class imbalance, or limited labeled graphs.
  • Issue 5. Inconsistent Evaluation Pipelines. The lack of a standardized evaluation pipeline results in inconsistent comparisons. Different works often use varying data splits, tuning protocols, and evaluation metrics, hindering fair assessment of model performance. A unified evaluation framework is needed for transparent and reliable comparisons.
To address these five issues, we introduce OpenGLT, a comprehensive framework designed to provide a fair and thorough assessment of GNNs for graph-level tasks. We begin by systematically categorizing existing GNNs for graph-level tasks into five distinct types and conduct an in-depth analysis of each type to understand their unique strengths and limitations. Next, we introduce a unified evaluation pipeline with standardized data splitting and experimental settings. This pipeline incorporates graph datasets from diverse domains, including biology, chemistry, social networks, and motif graphs, ensuring broad and representative evaluations. Furthermore, we comprehensively evaluate the effectiveness and efficiency of existing graph-level GNNs on both graph classification and graph regression tasks, such as cycle counting. Beyond standard tasks, we also evaluate GNNs in real-world scenarios, including noisy graphs, imbalanced datasets, and few-shot learning settings. This unified framework ensures a holistic and equitable evaluation of GNNs, promoting fair comparisons and advancing their applicability across various graph-level tasks.
We summarize the contributions of this paper on graph-level benchmarking as follows.
  • We systematically revisit GNNs for graph-level tasks and categorize them into five types, with in-depth analysis in Section 3.
  • We propose a unified open-source evaluation framework, OpenGLT, which covers diverse tasks, datasets, and scenarios.
  • We conduct extensive experiments on 13 graph classification and 13 graph regression datasets, systematically comparing the effectiveness and efficiency of state-of-the-art GNNs, providing insights into their relative strengths, limitations, and directions.
2. Preliminaries and Related Work
In this section, we first introduce graph-level tasks and then review existing experimental studies related to these graph-level tasks. Important notations are summarized in Table 1.
2.1. Graph-level Tasks
Graphs are a fundamental data structure to model relations between nodes in real-world datasets. Consequently, graph-level tasks, including classification and regression, are widely applied across domains, such as recommendation systems in social networks (Cohen, 2016; Linghu et al., 2020), molecular property and activity prediction in chemistry (Chen et al., 2024; Cohen, 2016; Mandal et al., 2022; Niazi and Mariam, 2023; Gilmer et al., 2017), and protein analysis, motif identification, or gene expression prediction in biology (Zang et al., 2023; Otal et al., 2024). Formally, we denote a graph as $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$, where $V_{i}$, $\mathbf{A}_{i}\in\{0,1\}^{|V_{i}|\times|V_{i}|}$, $\mathbf{X}_{i}\in\mathbb{R}^{|V_{i}|\times d_{x}}$, denote nodes, adjacency matrix, and node features, respectively. In general, graph-level tasks are to learn a mapping from $G_{i}$ to a label (classification) or a continuous value (regression).
2.2. Related Works of Experimental Studies
As summarized in Table 2, existing benchmarks and surveys (Zhiyao et al., 2024; Li et al., 2024c; Hu et al., 2020a; Errica et al., 2020; Wang et al., 2024b) often face four key limitations (Zhiyao et al., 2024; Li et al., 2024c; Hu et al., 2020a; Errica et al., 2020; Wang et al., 2024b): (i) a lack of systematic taxonomy for graph-level tasks; (ii) a predominant focus on node-based GNNs (Dwivedi et al., 2023; Hu et al., 2020a; Morris et al., 2020), leaving diverse architectures partially absent with poor categorization; (iii) the neglect of realistic scenarios such as noise, imbalance, and few-shot settings, which hinders robustness assessment; and (iv) a lack of comprehensive efficiency metrics for usability evaluation.
Table 1. Summary on important notations.
Symbols Meanings
$G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ The graph $G_{i}$
$y_{i}$ Discrete label or continuous value of $G_{i}$
$N^{l}_{i}{(v)},N_{i}(v)$ The $l$-hop and 1-hop neighbors of $v$ in $G_{i}$
$f_{\theta}$ GNN model
$l,L$ GNN layer index and the total layer number
$\mathbf{h}^{(l)}_{i}(v)$ The $l$-th layer representation of $v$ in $G_{i}$
$m^{(l)}_{i}(v)$ The $l$-th layer message aggregated intermediate for node $v$
$e_{uv}$ Edge feature of edge $(u,v)$ in $G_{i}$
$\mathbf{h}_{i}(v)$ Final node representation of $v$ in $G_{i}$
$\mathbf{H}^{(l)}_{i}(V_{j})$ $l$-th layer representation of nodes $V_{j}$ in $G_{i}$
$\mathbf{H}_{i}(V_{j})$ Final representation of nodes $V_{j}$ in $G_{i}$
$\mathrm{READOUT}(\cdot)$ Graph-level aggregation function producing $\mathbf{h}_{i}$ from node/cluster embeddings
$\mathbf{h}_{i}$ Graph representation of $G_{i}$
${y}^{*}_{i}$ Prediction of GNN $f_{\theta}$ for $G_{i}$
$\mathbf{S}_{i}^{(l)}$ Cluster assignment matrix at $l$-th layer
$\hat{G}^{*}_{i}(V_{i},\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i})$ The reconstructed graph for $G_{i}$
$\tilde{G}_{i},\hat{G}_{i}$ Augmented positive views for $G_{i}$
$s({\mathbf{h}}_{i},{\mathbf{h}_{j}})$ Similarity score between $G_{i}$ and $G_{j}$
$\mathcal{L}_{task}(\cdot)$ Task loss (Equation (4))
$\mathcal{L}_{ssl}(\cdot)$ Self-supervised learning loss (Equation (11))
$\mathcal{L}_{cl}(\cdot)$ Contrastive learning loss (Equation (15))
Table 2. Summary of existing surveys and experimental studies on GNNs for graph-level tasks. Sur. and Exp. denote Survey and Experiments, respectively. Taxo. denotes Taxonomy, Subg. denotes Subgraph. Additionally, FewS., Imba., Effect., and Effic. denote Few-shot, Imbalanced, Effectiveness, and Efficiency, respectively.
Paper Paper Type Taxo. GNN Type Data Type Scenarios Eval Metric
Sur. Exp. Node Pool Subg. GL SSL SN BIO CHE MC Clean Noise FewS. Imba. Effect. Effic.
GNNS (Wu et al., 2020)
TUD (Morris et al., 2020)
OGB (Hu et al., 2020a)
GNNB (Dwivedi et al., 2023)
ReGCB (Li et al., 2024c)
FGNNB (Errica et al., 2020)
GPB (Wang et al., 2024b)
OpenGLT
3. Graph Neural Networks for Graph-level Tasks
Recently, GNNs (Demirci et al., 2022; Zou et al., 2023; Shao et al., 2022; Guliyev et al., 2024; Li and Chen, 2023; Huang et al., 2024; Wang et al., 2024a; Gao et al., 2024; Zhang et al., 2023) have emerged as powerful tools to learn node representations, due to their ability to capture complex relationships within graph structures. Existing GNNs for graph-level tasks can generally be categorized into five types: node-based, pooling-based, subgraph-based, graph learning-based, and self-supervised learning-based GNNs.
3.1. Node-based GNNs
As shown in Figure 1 (a), node-based GNNs first learn latent representations for individual nodes, which then are used to form a graph-level representation. Generally, each GNN $f_{\theta}$ consists of two fundamental operations, i.e., AGG($\cdot$) and COM($\cdot$(Hamilton et al., 2017; Li and Chen, 2021), which can be parameterized by learnable matrices $\mathbf{W}_{agg}$ and $\mathbf{W}_{com}$, respectively. Formally, given a graph $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$ and a node $v\in V_{i}$, in the $l$-th layer, $\mathsf{AGG}^{(l)}(\cdot)$ in Equation (1) first aggregates the hidden representation $\mathbf{h}_{i}^{(l-1)}(u)$ and edge representation $\mathbf{e}_{uv}$ of each neighbor $u$ of node $v$ and obtain the aggregated result $\mathbf{m}^{(l)}_{i}(v)$. Subsequently, as shown in Equation (2), $\mathsf{COM}^{(l)}(\cdot)$ combines the aggregated neighbor information $\mathbf{m}_{i}^{(l)}(v)$ and the latest information $\mathbf{h}_{i}^{(l-1)}(v)$ of node $v$ to obtain the $l$-th layer graph-level representation $\mathbf{h}_{i}^{(l)}(v)$.
(1) $\mathbf{m}_{i}^{(l)}(v)$ $=\mathsf{AGG}^{(l)}\left(\{(\mathbf{h}_{i}^{(l-1)}(u),\mathbf{e}_{uv}):u\in\mathcal{N}_{i}(v)\}\right)$
(2) $\mathbf{h}_{i}^{(l)}(v)$ $=\sigma\left(\mathsf{COM}^{(l)}\left(\mathbf{h}_{i}^{(l-1)}(v),\mathbf{m}_{i}^{(l)}(v)\right)\right),$
where $\sigma$ denotes a non-linear function (e.g., ReLU (Li and Yuan, 2017)). $\mathbf{h}_{i}^{(0)}(v)$ is initialized as $\mathbf{X}_{i}[v]$. After $L$ layers aggregation, we can obtain node representations $\mathbf{H}_{i}(V_{i})\in\mathbb{R}^{|V_{i}|\times d_{H}}$. Then, the graph representation $\mathbf{h}_{i}$ of $G_{i}$ can be summarized from the node representation $\mathbf{H}_{i}(V_{i})$ by a $\mathsf{READOUT}$ function (Li et al., 2024a) as follows:
(3) $\mathbf{h}_{i}=\mathsf{READOUT}(\mathbf{H}_{i}(V_{i})),$
where READOUT is a permutation-invariant function, such as AVERAGE, SUM, or MAX functions (Li et al., 2024a). The $\mathsf{SUM}$ function as an example is $\mathbf{h}_{i}=\sum_{v\in V_{i}}\mathbf{H}_{i}(V_{i})[v]$.
Model Optimization. Then, we can use a decoder (e.g., 2-layer Multi-Layer Perceptrons (MLPs) (Zhu and others, 2021)) to predict discrete class or continuous property value for each graph $G_{i}$ based $\mathbf{h}_{i}$. In summary, given the labeled training dataset $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$ and the prediction $y^{*}_{i}$ for each graph $G_{i}$ by the GNN $f_{\theta}$, the GNN $f_{\theta}$ can be optimized by minimizing the task loss as follows:
(4) $\theta^{*}=\arg\min_{\theta}\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{\mathcal{L}_{task}(f_{\theta},G_{i},y_{i})}.$
(5) $s.t.\quad$ $\mathcal{L}_{task}(y^{*}_{i},y_{i})=\begin{cases}-\sum_{y\in{y}_{i}}y\log y^{*}_{i}[y],&{y}_{i}\in\{0,1\}^{|\mathcal{Y}|}\\ \|y^{*}_{i}-y_{i}\|_{2},&y_{i}\in\mathbb{R}\end{cases}$
3.2. Hierarchical Pooling-based GNNs
As shown in Figure 1 (b), hierarchical pooling (HP)-based GNNs are designed to capture hierarchical structures in the graph, which can reduce the graph size progressively while preserving its essential structural information. This is achieved by clustering nodes into groups and summarizing their features, allowing the model to focus on condensed graph structures and finally obtain the representation of each graph.
Figure 1. The five types of current GNNs for graph-level tasks.
As mentioned in Equations (1) and (2), at the $l$-th layer, node-based GNNs learn node representations using the adjacency matrix $\mathbf{A}_{i}$ and the node hidden representations $\mathbf{H}_{i}^{(l-1)}(V_{i})$. In contrast, at the $l$-th layer, HP-based GNNs first generate a new cluster assignment matrix $\mathbf{S}_{i}^{(l)}\in\{0,1\}^{n_{l-1}\times n_{l}}$, where $n_{l}<n_{l-1}$. This matrix maps the nodes $V_{i}^{(l-1)}$ from the (l1)(l-1)-th layer to a cluster index in $\{1,\ldots,n_{l}\}$. The new adjacency matrix $\mathbf{A}_{i}^{(l)}\in\mathbb{R}^{n_{l}\times n_{l}}$ is then reconstructed based on the cluster assignment matrix $\mathbf{S}_{i}^{(l)}$ as follows:
(6) $\mathbf{A}_{i}^{(l)}=\mathbf{S}_{i}^{(l)\top}\mathbf{A}_{i}^{(l-1)}\mathbf{S}_{i}^{(l)}.$
Subsequently, $n_{l}$ new nodes $V_{i}^{(l)}$ are created to represent the $n_{l}$ clusters. Then, the node hidden representations $\mathbf{H}^{(l-1)}_{i}(V^{(l)}_{i})$ can obtained by a READOUT operation on nodes each cluster $k\in\{1,\ldots,n_{l}\}$ as follows:
(7) $\mathbf{H}^{(l-1)}_{i}(V^{(l)}_{i})[{k}]=\textsf{READOUT}(\ \mathbf{H}^{(l-1)}_{i}(V^{(l-1)}_{i})[V_{k}]),$
where $V_{k}=\{u\mid\mathbf{S}_{i}^{(l)}[u][k]=1\}$ are nodes in the cluster $k$.
Depending on the technique of generating the cluster matrix $\mathbf{S}_{i}^{(l)}$, existing pooling-based GNNs can be classified into three types.
  • Similarity-based. Similarity-based approaches (Mesquita et al., 2020; Dhillon et al., 2007; Li et al., 2022b) cluster nodes based on the similarity of nodes features or structures. These methods rely on predefined similarity metrics such as cosine similarity of node features, or graph partitioning algorithms to generate the cluster assignment matrix. Graclus (Mesquita et al., 2020; Dhillon et al., 2007) and CC-GNN (Li et al., 2022b) assign nodes to different clusters based on feature similarity and graph structure.
  • Node Dropping-based. Node dropping-based methods (Cangea et al., 2018; Gao and Ji, 2019; Lee et al., 2019) focus on reducing the number of nodes in the graph by learning an importance score for each node. Nodes with the highest scores are retained, while others are dropped, effectively shrinking the graph size in a controlled manner. TopKPool (Cangea et al., 2018; Gao and Ji, 2019) and SAGPool (Lee et al., 2019) aim to learn an importance score for each node and retain only the top-$n_{l}$ nodes with the highest scores at layer $l$. In other words, they only assign one node for each cluster and directly drop the other nodes.
  • Learning-based. Learning-based approaches (Ying et al., 2018; Bianchi et al., 2020; Vaswani, 2017; Baek et al., 2021; Diehl, 2019) use neural network architectures to learn the cluster assignment matrix $\mathbf{S}_{i}^{(l)}$ based on the graph’s node features and structure, which adaptively learn how to cluster nodes in a way that preserves the graph’s essential information. Specifically, DiffPool (Ying et al., 2018) and MinCutPool (Bianchi et al., 2020) use non-linear neural networks and and GMT (Baek et al., 2021) use the multi-head attention mechanism (Vaswani, 2017), to learn a cluster matrix $\mathbf{S}_{i}^{(l)}$ based on the learned node features $\mathbf{H}_{i}^{(l-1)}$ and adjacent matrix $\mathbf{A}_{i}^{(l-1)}$. EdgePool (Diehl, 2019) proposes to learn the edge score between each pair of connected nodes and then generate the cluster matrix accordingly.
3.3. Subgraph-based GNNs
Recent studies propose subgraph-based GNNs with enhanced expressive power, enabling them to more effectively capture substructure information. As shown in Figure 1 (c), the core idea is to decompose the input graph into a set of overlapping or non-overlapping subgraphs and then learn representations for each subgraph to improve the final graph representation. Formally, given a graph $G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})$, existing models first generate $n_{s}$ subgraphs ${G_{i,j}(V_{i,j},\mathbf{A}_{i,j},\mathbf{X}_{i,j})}_{j=1}^{n_{s}}$. Then, for each subgraph $G_{i,j}$, the node representation of $v\in V_{i,j}$ is learned by Equations (1) and (2). In this approach, each node $v\in V_{i}$ can have multiple representations because it may appear in multiple sampled subgraphs. To derive the final node representation $\mathbf{h}_{i}(v)$, these representations are merged. A common approach is to use a READOUT function to combine them, as follows:
(8) $\mathbf{h}_{i}(v)=\textsf{READOUT}(\mathbf{h}_{i,j}(v)|v\in V_{i,j}).$
Depending on the techniques of generating the subgraphs, current subgraph-based approaches can be classified into three types.
  • Graph Element Deletion-based. These approaches  (Cotta et al., 2021; Ding et al., 2024; Papp et al., 2021; Bevilacqua et al., 2022) propose deleting specific nodes or edges to create subgraphs, enabling GNNs to focus on the most informative parts of the graph by removing noise or less relevant elements. For example, DropGNN (Papp et al., 2021) and ESAN (Bevilacqua et al., 2022) generate subgraphs through random edge deletion to enhance model expressiveness. Conversely, SGOOD (Ding et al., 2024) abstracts a superstructure from the original graphs and applies sampling and edge deletion on this superstructure, thereby creating a more diverse graphs.
  • Rooted Subgraph-based. These approaches (Bevilacqua et al., 2024; You et al., 2021; Yan et al., 2024; Frasca et al., 2022; Huang et al., 2022; Papp and Wattenhofer, 2022; Qian et al., 2022; Yang et al., 2023; Zhang and Li, 2021; Zhao et al., 2022b) focus on generating subgraphs centered around specific nodes, referred to as root nodes, with the aim of capturing the structural role and local topology of the root node within the subgraph. The primary motivation is to enhance the representational power of GNNs by encoding the root node’s position within the subgraph, as well as its relationships to other nodes within the same subgraph. Specifically, I2GNN (Huang et al., 2022), ECS (Yan et al., 2024), and ID-GNN (You et al., 2021) propose to append side information to the root nodes or edges to capture the position information, such as an ID identifier for root nodes (e.g., the root $v$ is 1 and other nodes are 0) (Huang et al., 2022; You et al., 2021), and node degree and shortest distance information (Yan et al., 2024). Also, NestGNN (Zhang and Li, 2021) and GNN-AK (Zhao et al., 2022b) use rooted subgraphs with different hops to capture the hierarchical relationships within graphs for root nodes.
  • $k$-hop Subgraph-based. Similar to rooted subgraph-based approaches (Nikolentzos et al., 2020; Feng et al., 2022; Abu-El-Haija et al., 2019; Yao et al., 2023; Sandfelder et al., 2021; Wang et al., 2021), $k$-hop subgraph-based approaches construct subgraphs based on the $k$-hop neighborhood of each node $v$, focusing on capturing the local structure around each node. Unlike rooted subgraph-based approaches, which aggregate information from 1-hop neighbors, $k$-hop subgraph-based approaches aggregate information not only from 1-hop neighbors but also directly from nodes up to $k$ hops away. Specifically, MixHop (Abu-El-Haija et al., 2019) utilizes a graph diffusion kernel to gather multi-hop neighbors and computes the final representation. SEK-GNN (Yao et al., 2023), KP-GNN (Feng et al., 2022), EGO-GNN (Sandfelder et al., 2021), and $k$-hop GNN (Nikolentzos et al., 2020) progressively updates node representations by aggregating information from neighbors within $k$-hops. MAGNA (Wang et al., 2021) proposes to learn more accurate weights between each pair of nodes based on all the paths between them within $k$-hops.
3.4. Graph Learning-based GNNs
Due to uncertainty and complexity in data collection, real-world graphs often contain redundant, biased, or noisy edges and features. When operating on such imperfect structures, vanilla GNNs may learn spurious correlations and thus fail to produce reliable graph representations, ultimately leading to incorrect predictions. To mitigate this issue, as shown in Figure 1 (d), recent work (Fatemi et al., 2023; Zhiyao et al., 2024; Li et al., 2024d) propose to learn from purified or reconstructed graph structure and enhanced node features that better reflect the underlying signal to improve the quality of the learned graph representations. Given a labeled graph set $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{n}$, the graph learning-based approaches can be formulated as bi-level optimization problem.
(9) $\theta^{*}=\min_{\theta\in\Theta}{\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\mathcal{L}_{task}({f}_{\theta},\hat{G}^{*}_{i}(V_{i},\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i}),y_{i})}.$
$s.t.\quad$ $\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i}=\arg\min_{\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}}\mathcal{L}_{gl}\bigl(f_{\theta^{*}},\hat{G}_{i}(V_{i},\hat{\mathbf{A}}_{i},\hat{\mathbf{X}}_{i}),y_{i}\bigr),$
(10) $\hfill\forall G_{i}\in\mathcal{G}$
At the low level in Equation (10), current approaches propose different graph learning objectives $\mathcal{L}_{gl}(\cdot)$ to reconstruct graph structure $\mathbf{A}_{i}$ and node features $\mathbf{X}_{i}$. Then, in Equation (9), $\hat{G}^{*}_{i}(V_{i},\hat{\mathbf{A}}^{*}_{i},\hat{\mathbf{X}}^{*}_{i})$ will be used to optimize the GNNs by the loss function in Equation (4).
Depending on the techniques of reconstructing graphs, current GL-based GNNs can be categorized into three types.
  • Preprocessing-based. Preprocessing-based approa-
    ches (Wu et al., 2019; Li et al., 2022a; Entezari et al., 2020) reconstruct graphs and then use them to optimize GNNs directly. The basic idea is to first reveal the common graph patterns and modify the graph structure and node features to recover these patterns. For example, GNN-Jaccard (Wu et al., 2019) and GNAT (Li et al., 2022a) observe that similar nodes tend to be connected by edges. Therefore, they remove edges between dissimilar nodes and add edges between similar ones. Also, GNN-SVD (Entezari et al., 2020) observes that noisy edges and node features tend to increase the rank of the adjacency matrix. Therefore, GNN-SVD reconstructs graphs by decreasing the rank of the adjacency matrix.
  • Jointly Training-based. In contrast to static preprocessing-based methods, jointly training-based approaches aim to optimize the GNN while iteratively reconstructing the graph structure and node features alongside the GNN model parameters. This iterative bi-level optimization enables the GNN to dynamically adapt the graph structure and features based on task-specific objectives. For instance, several approaches (Jin et al., 2021, 2020b; Li et al., 2024b; Franceschi et al., 2019; Sun et al., 2022; Luo et al., 2021), such as ADGNN (Li et al., 2024b), ProGNN (Jin et al., 2020b), and SimPGCN (Jin et al., 2021), propose reconstructing the edges between each pair of nodes by minimizing the GNN loss on the reconstructed structure while also reducing the rank of the adjacency matrix. Alternatively, several approaches (Zhang et al., 2019; Zhou et al., 2024; Wang et al., 2023a), such as MOSGSL (Zhou et al., 2024) and HGP-SL (Zhang et al., 2019), focus on first partitioning each graph into subgraphs based on node similarities and predefined motifs, and then reconstructing the edges between these subgraphs rather than at the individual node level.
3.5. Self Supervised Learning-based GNNs
Self-supervised learning (SSL) has become a powerful paradigm to pretrain GNNs without the need for labeled data, which can capture the node patterns and graph patterns. As shown in Figure 1 (e), the key idea of SSL approaches is to create supervised signals directly from the structure and node features of the unlabeled graph itself, leveraging the graph’s inherent properties to guide the learning process. Formally, given a set of unlabeled graphs $\mathcal{UG}=\{G_{i}(V_{i},\mathbf{A}_{i},\mathbf{X}_{i})\}_{i=1}^{|\mathcal{UG}|}$, the GNN $f_{\theta}$ is pretrained as follows:
(11) $\theta^{\prime}=\arg\min_{\theta}\frac{1}{|\mathcal{UG}|}\sum_{G_{i}\in\mathcal{UG}}\mathcal{L}_{ssl}(f_{\theta},G_{i},Signal_{i}),$
where $Signal_{i}$ is the supervised signals from the unlabeled graph $G_{i}$ and $\theta^{\prime}$ is the optimized GNN parameters. Then, the pretrained $f_{\theta^{\prime}}$ can be used to predict graph labels or properties. Formally, given the set of labeled graphs $\mathcal{LG}=\{G_{j}(V_{j}\mathbf{A}_{j},\mathbf{X}_{j}),y_{j}\}_{j=1}^{|\mathcal{LG}|}$, the GNN $f_{\theta^{\prime}}$ is optimized as follows:
(12) $\theta^{*}=\arg\min_{\theta^{\prime}}\frac{1}{|\mathcal{LG}|}\sum_{G_{j}\in\mathcal{LG}}\mathcal{L}_{task}(f_{\theta^{\prime}},G_{j},y_{j}),$
where the task loss $\mathcal{L}_{task}(\cdot)$ is defined in Equation (4).
Depending on the specific technique used to auto-generate supervised signals from unlabeled graphs, SSL-based GNN approaches can be broadly categorized into two main paradigms.
  • Pretext Task-based. Pretext task-based approaches  (Hou et al., 2022; Inae et al., 2023; Zhang et al., 2021; Zang et al., 2023; Hu et al., 2020b; Jin et al., 2020a; Wang et al., 2022b) design auxiliary tasks to help the model learn useful representations from the graph structure and features without requiring external labels. Current approaches design diverse node-level and graph-level tasks to capture intrinsic graph patterns (Wang et al., 2022b), such as predicting node attribute, node degree, and the node number. For example, HMGNN (Zang et al., 2023) pretrains GNNs by predicting the links between nodes and the number of nodes in each graph. MGSSL (Zhang et al., 2021) first masks the edges among motifs and then predicts the missing edges. MoAMa (Inae et al., 2023) first masks node features and then uses GNNs to reconstruct the node features based on the masked graphs. GraphMAE (Hou et al., 2022) and GPTGNN (Hu et al., 2020b) predict both node attributes and edges.
  • Graph Contrastive Learning-based. Graph contrastive learning (GCL)-based GNNs (Wang et al., 2022a; Perozzi and others, 2014; Lee et al., 2022; Hassani and Khasahmadi, 2020; Yuan et al., 2021; Sun et al., 2024) aim to learn representations by contrasting positive and negative examples. The basic idea is to maximize the similarity between different views or augmentations of the same node or graph (positive pairs) while minimizing the similarity between different nodes or graphs (negative pairs). In such a way, GCL-based approaches can distinguish the similarity between nodes and graphs. In general, the SSL loss $\mathcal{L}_{ssl}(\cdot)$ in Equation (11) can be formulated as follows.
    (13) $\theta^{\prime}=\arg\min_{\theta}{\frac{1}{|\mathcal{UG}|}\sum_{G_{i}\in\mathcal{UG}}\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})}.$
    (14) $s.t.\$$\tilde{G}_{i},\hat{G}_{i}=\arg\min_{\tilde{G}_{i},\hat{G}_{i}}{\mathcal{L}_{positive}(G_{i},\mathcal{T})},\forall G_{i}\in\mathcal{UG},$
    where $\mathcal{L}_{cl}(\cdot)$ is the contrastive loss. Additionally, $\mathcal{L}_{positive}(\cdot)$ denotes the view generation loss used to generate two positive views (i.e., $\tilde{G}_{i}$ and $\hat{G}_{i}$ for each graph $G_{i}$) where $\mathcal{T}$ is a set of augmentation operations for generating the views. Negative views ($Neg_{i}$) are a set of graphs, which are typically randomly sampled from all graphs or from graphs whose labels differ from $G_{i}$. One typical of the contrastive loss $\mathcal{L}_{cl}(f_{\theta},\hat{G}_{i},\tilde{G}_{i},Neg_{i})$ in Equation (13) can be defined based on InfoNCE loss (Zhu and others, 2021; Yeh et al., 2022) as follows:
    (15) $\mathcal{L}_{cl}(\cdot)=-\log\frac{\text{s}(\mathbf{\hat{h}}_{i},\mathbf{\tilde{h}}_{i})}{\sum_{G^{\prime}_{i}\in\{\hat{G}_{i},\tilde{G}_{i}\}}\sum_{G_{j}\in A(G_{i})\}}s(\mathbf{\hat{h}}^{\prime}_{i},\mathbf{\tilde{h}}_{j})},$
    where $A(G_{i})=\hat{G}_{i}\cup\tilde{G}_{i}\cup{Neg}_{i}$ includes positive views and negative samples of $G_{i}$, and $\hat{\mathbf{h}}_{i}$ and $\tilde{\mathbf{h}}_{i}$ are the representations of graph $\hat{G}_{i}$ and $\tilde{G}_{i}$, respectively. Also, $s({\mathbf{h}}_{i},{\mathbf{h}_{j}})=\exp(cosine({\mathbf{h}}_{i},{\mathbf{h}}_{j})/\tau)$ is the similarity score between ${\mathbf{h}}_{i}$ and ${\mathbf{h}}_{j}$ based the parameter $\tau$.
    The techniques in Equation (14) are employed to generate positive views for each node. Specifically, similarity-based methods (Wang et al., 2022a; Perozzi and others, 2014; Lee et al., 2022) utilize both node features and graph structures to identify nodes in the original graph, which are similar to a given target node, as the positive nodes. Also, diffusion-based approaches (Hassani and Khasahmadi, 2020; Yuan et al., 2021; Sun et al., 2024) modify the graph’s structure (e.g., by deleting or adding edges) based on global topological information, such as personalized PageRank (PPR) (Haveliwala, 1999) and motif-preserving (Sun et al., 2024). Perturbation-based methods (Thakoor et al., 2021; Zhu et al., 2020; Zhu and others, 2021; Li et al., 2024a), on the other hand, introduce changes to both the graph’s structure and node features.
4. Experiments
4.1. Evaluation Framework OpenGLT
As shown in Figure 2, we design a flexible and end-to-end evaluation framework OpenGLT that consists of three key components: the data-level, model-level, and evaluation-level components. It ensures fair and comprehensive comparisons across different methods.
  • Data-level Component. We collected 13 datasets for graph classification spanning diverse domains, including social networks, chemical graphs, and molecular graphs, as well as 13 datasets for graph regression like cycle and path counting. Additionally, we provide a unified data processing pipeline, which includes data transformation and standardized data splits tailored to various scenarios, such as imbalanced and few-shot settings. All datasets are formatted into a unified loader to facilitate fair comparisons across different methods and support future extensions.
  • Model-level Component. At the model level, we provide more than 20 GNN models to support comprehensive evaluations. These include node-based models, pooling-based models, subgraph-based models, graph-level models (GL-based), and self-supervised learning-based models (SSL-based). Additionally, we incorporate various optimization strategies tailored to GNNs, such as neighbor sampling, subgraph sampling, and graph partitioning, to enhance scalability and performance.
  • Evaluation-level Component. It focuses on ensuring fair and comprehensive comparisons of model performance. We utilize a wide range of metrics, including Macro-F1, Micro-F1, Mean Absolute Error (MAE), and $R^{2}$ scores, as well as resource-based measures such as runtime and memory consumption. To aid in the interpretation of results, we provide multiple visualization options, including tables, column charts, and line charts, enabling clear and insightful comparisons across different methods.
Our framework is designed to be user-friendly and highly extensible. Users can easily integrate new GNN models by following a standardized interface and registering them in the model library. Similarly, new datasets can be incorporated by formatting them into our unified loader, which handles preprocessing and splits automatically. The evaluation module also supports the addition of custom metrics and visualizations. To assist users, we provide comprehensive documentation with examples, ensuring that the framework can be readily extended to meet diverse research needs.
Figure 2. Evaluation framework.
4.2. Datasets
We evaluate GNNs on four domains with 26 datasets: social networks (SN), biology (BIO), chemistry (CHE), and motif counting (MC). For dataset splitting, if a standard split is publicly available and commonly used in existing works, we adopt the standard split. If there is no standard split, we use a 10-fold cross-validation approach. Detailed statistics and description are in Table 3.
4.2.1. Social Networks
  • IMDB-BINARY (I-B) (Morris et al., 2020) and IMDB-MULTI (I-M) (Morris et al., 2020). Two movie datasets from IMDB (IMDb, 2024) where each graph represents a movie, with nodes as actors and edges indicating co-appearance. The task of IMDB-B is to classify movies into two genres: Action and Romance. IMDB-M extends this to a three-class setting: Comedy, Action, and Drama.
  • REDDIT (RED) (Morris et al., 2020). We use the binary version (REDDIT-BINARY) of a dataset from Reddit (Reddit, 2024), where each graph represents a discussion thread. Nodes correspond to users, and edges indicate interactions. The task is to classify threads into two categories: Q&A-based and Discussion-based.
  • COLLAB (COL) (Morris et al., 2020). A scientific collaboration dataset where each graph represents an ego-network of a researcher, with nodes as researchers and edges indicating co-authorship. The task is to classify the networks into one of three research fields: High Energy Physics, Condensed Matter Physics, or Astrophysics.
4.2.2. Biology
  • PROTEINS (PRO) (Morris et al., 2020) and DD (Morris et al., 2020). Two binary classification datasets where each graph represents a protein. The task is to classify a protein molecule as either an enzyme or a non-enzyme. In PROTEINS, nodes represent secondary structure elements (SSEs), while in DD, nodes are amino acids; edges in both indicate spatial proximity.
  • ENZYMES (ENZ) (Morris et al., 2020). A six-class classification dataset from the BRENDA enzyme database. Each graph represents a protein, in which nodes represent SSEs with features encoding atomic properties, and edges denote spatial proximity. The task is to classify proteins into one of six enzyme commission (EC) classes.
  • MolHIV (HIV) (Hu et al., 2020a) and MolTox21 (TOX) (Hu et al., 2020a). Two molecular property prediction datasets where nodes represent atoms and edges denote chemical bonds. MolHIV is a single-label binary dataset and the task is to predict whether a molecule inhibits HIV replication. MolTox21 is a 12-label binary dataset, each label represents a specific toxicological target.
4.2.3. Chemistry
MUTAG, NCI1, MolBACE, and MolPCBA are molecular graph datasets, where nodes represent atoms and edges represent chemical bonds.
  • MUTAG (MUT) (Morris et al., 2020) and NCI1 (NCI) (Morris et al., 2020). In both datasets, node features are one-hot encodings of atomic species (e.g., C, N, O). They are binary classification tasks: MUTAG labels indicate whether a molecule is mutagenic, while NCI1 labels indicate anti-cancer activity in cell-based assays.
  • MolBACE (BAC) (Hu et al., 2020a) and MolPCBA (PCB) (Hu et al., 2020a). In both datasets, node features are derived from RDKit, including properties such as chirality and aromaticity. MolBACE is a single-label binary task for predicting whether a molecule inhibits the BACE-1 enzyme (linked to Alzheimer’s disease). MolPCBA is a 128-label binary task, each representing activity in a specific biological assay.
4.2.4. Motif-Counting Graphs
The graphs and their splits are sourced from (Chen et al., 2020) and have been widely used in follow-up works (Chen et al., 2020; Zhao et al., 2022b; Huang et al., 2022; Yan et al., 2024). Each graph is labeled with the counts of specific substructures, including Cycles (3-8 nodes), Paths (4-6 nodes), 4-Cliques, Tailed Triangles, Chordal Cycles, and Triangle-Rectangles. Graphs are generated by sampling a $d$-regular graph with $m$ nodes and randomly deleting $m$ edges, where $(m,d)$ is uniformly chosen from $\{(10,6),(15,6),(20,5),(30,5)\}$. The ground-truth counts are computed using the subgraph counting algorithm proposed in (Shervashidze et al., 2009).
Table 3. Data statistics. $\overline{\mbox{Nodes}}$ and $\overline{\mbox{Edges}}$ denote the average number of nodes and edges per graph, respectively. Split is the data partition strategy, where 10-fold denotes 10-fold cross-validation, others denote train/validation/test splits.
Type Dataset Graphs $\overline{\textbf{Nodes}}$ $\overline{\textbf{Edges}}$ Classes Split
SN IMDB-B (I-B) 1,000 19.8 96.5 2 10-fold
IMDB-M (I-M) 1,500 13.0 65.9 3 10-fold
REDDIT-B (RED) 2,000 429.6 497.8 2 10-fold
COLLAB (COL) 5,000 74.5 2457.8 3 10-fold
BIO PROTEINS (PRO) 1,113 39.1 72.8 2 10-fold
$D$ 1,178 284.3 715.7 2 10-fold
ENZYMES (ENZ) 600 32.6 62.1 6 10-fold
MolHIV (HIV) 41,127 25.5 54.9 2 8/1/1
MolTox21 (TOX) 7,831 18.9 39.2 12 8/1/1
CHE MUTAG (MUT) 188 17.9 19.8 2 10-fold
NCI1 (NCI) 4,110 29.9 32.3 2 10-fold
MolBACE (BAC) 1,513 34.1 36.9 2 8/1/1
MolPCBA (PCB) 437,929 26.0 28.1 128 8/1/1
MC {3,4,5,6,7,8}-Cycle 5,000 18.8 31.3 1 3/2/5
{4,5,6}-Path 5,000 18.8 31.3 1 3/2/5
4-Clique 5,000 18.8 31.3 1 3/2/5
Tailed Tri. 5,000 18.8 31.3 1 3/2/5
Chor. Cyc. 5,000 18.8 31.3 1 3/2/5
Tri. Rec. 5,000 18.8 31.3 1 3/2/5
4.3. Evaluated GNNs
We comprehensively evaluate 16 representative and effective GNNs across five categories as follows.
4.3.1. Node-based GNNs
  • GCN (Kipf and Welling, 2016). As a foundational GNN, GCN learns node features by aggregating information from their neighborhood, effectively performing a weighted average of neighbor representations in each layer.
  • GIN (Xu et al., 2019). GIN utilizes a Multi-Layer Perceptron (MLP) to update node representations based on a sum of its neighbors, theoretically matching the capability of the Weisfeiler-Lehman (1-WL) graph isomorphism test.
  • GraphSAGE (SAGE) (Hamilton et al., 2017). SAGE introduces a sampling-based neighborhood aggregation strategy. For each node, it uniformly samples a fixed number of neighbors to generate embeddings, making the learning process scalable and inductive.
  • PNA (Corso et al., 2020). PNA combines multiple aggregators (e.g., mean, max, sum) with degree-based scalers to capture the distribution of neighbor features, allowing the model to distinguish diverse neighbor patterns better than single-aggregator methods.
4.3.2. Pooling-based GNNs
  • TopK (Cangea et al., 2018; Gao and Ji, 2019). TopK uses a trainable projection vector to score nodes, with the top-k nodes passing GNN backbones selected to form the graph for the subsequent layer.
  • GMT (Baek et al., 2021). GMT introduces a hierarchical pooling method where nodes are adaptively grouped into clusters using a transformer-based attention mechanism, allowing the model to learn soft cluster assignments and generate coarse-grained graph representations at multiple scales.
  • EdgePool (EPool) (Diehl, 2019). EPool learns a hierarchical representation by iteratively contracting edges. It scores edges based on their importance for connecting nodes and merges the corresponding node pairs, effectively simplifying the graph topology layer by layer.
4.3.3. Subgraph-based GNNs
  • ECS (Yan et al., 2024). ECS boosts the GNN backones by injecting pre-calculated structural information. It computes embeddings that encode subgraph-centric distances, allowing the model to count substructures with provable guarantees.
  • GNNAK+ (AK+) (Zhao et al., 2022a). AK+ shifts the computational unit from single-node neighborhoods to k-hop subgraphs, thereby enabling the GNN backbones to capture more complex, high-order structural patterns.
  • I2GNN (I2) (Huang et al., 2022). I2GNN breaks structural symmetries by assigning unique identifiers to the central root node and its neighbors within each sampled subgraph, enabling finer-grained representation learning.
4.3.4. GL-based GNNs
  • VIBGSL (VIB) (Sun et al., 2022). VIB uses the Information Bottleneck principle to mask irrelevant features and learn a task-relevant graph structure for better modeling the key patterns.
  • HGP-SL (HGP) (Zhang et al., 2019). HGP selects key nodes to form an induced subgraph and refines the graph structure using sparse attention-based structure learning.
  • MOSGSL (MO) (Zhou et al., 2024). MO reconstructs key subgraphs using a motif-driven guidance module that dynamically captures and aligns discriminative structural patterns.
4.3.5. SSL-based GNNs
  • RGC (Sun et al., 2024). RGC employs a diverse-curvature GCN and motif-aware contrastive learning on a stable kernel layer to robustly model motif regularity and learn node representations.
  • MVGRL (MVG) (Hassani and Khasahmadi, 2020). MVG learns graph representations by contrasting node embeddings from the original and diffusion-based views of the graph.
  • GCA (Zhu and others, 2021). GCA learns graph representations by contrasting node embeddings across adaptively augmented views that preserve important structural and semantic information.
4.4. Evaluation Metric
We evaluate the performance of GNNs using effectiveness and efficiency metrics as follows.
4.4.1. Effectiveness Metric
Given a graph set $\mathcal{LG}=\{(G_{i},y_{i})\}_{i=1}^{|\mathcal{LG}|}$, we denote the prediction label of each graph $G_{i}$ as $\hat{y}_{i}$. For graph classification tasks, we use the Strict Accuracy (Acc), Micro-F1 (Mi-F1), and Macro-F1 (Ma-F1). Particularly, if each graph only has one label, the Micro-F1 is same as Accuracy.
  • Strict Accuracy (Acc). Strict accuracy is used to measure the proportion of exact matches between predicted and true labels. Strict accuracy is defined as $Acc=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{\mathbb{I}(y_{i}=\hat{y}_{i})}$, where $\mathbb{I}(y_{i}=\hat{y}_{i})=1$ if only $y_{i}=\hat{y}_{i}$.
  • Micro-F1 (Mi-F1). Micro-F1 is a performance metric that considers the overall precision and recall across all instances in the dataset. The Micro-precision is defined as $Mi\text{-}P=\frac{\sum_{G_{i}\in\mathcal{LG}}{|y_{i}\cap\hat{y}_{i}|}}{\sum_{G_{j}\in\mathcal{LG}}{|\hat{y}_{j}|}}$ and Micro-recall is $Mi\text{-}R=\frac{\sum_{G_{i}\in\mathcal{LG}}{|y_{i}\cap\hat{y}_{i}|}}{\sum_{G_{j}\in\mathcal{LG}}{|{y}_{j}|}}$. Then, the Micro-F1 is defined as $Mi\textit{-}F1=\frac{2\times Mi\text{-}P\times Mi\text{-}R}{Mi\text{-}P+Mi\text{-}R}$.
  • Macro-F1 (Ma-F1). Macro-F1 evaluates the average performance of precision and recall across all instances, treating each equally regardless of size. The Macro-precision is defined as $Ma\text{-}P=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|\hat{y}_{i}|}$ and Macro-recall is defined as $Ma\text{-}R=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}\frac{|y_{i}\cap\hat{y}_{i}|}{|y_{i}|}$, so the Macro-F1 is defined as $Ma\text{-}F1=\frac{2\times Ma\text{-}P\times Ma\text{-}R}{Ma\text{-}P+Ma\text{-}R}$.
For graph regression tasks, we use the Mean Absolute Error (MAE) and R2 as follows.
  • Mean Absolute Error (MAE). Mean Absolute Error measures the average magnitude of errors between predicted and true values. MAE is defined as $MAE=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}{|y_{i}-\hat{y}_{i}|}$. Lower $MAE$ indicates better performance.
  • R2. R2 evaluates the proportion of variance in the true values that is captured by the predicted values. The R2 is defined as $R2=1-\frac{\sum_{G_{i}\in\mathcal{LG}}(y_{i}-\hat{y}_{i})^{2}}{\sum_{G_{i}\in\mathcal{LG}}(y_{i}-\bar{y})^{2}}\in[0,1]$, where $\bar{y}=\frac{1}{|\mathcal{LG}|}\sum_{G_{i}\in\mathcal{LG}}y_{i}$ is the mean of the true values. Higher $R2$ indicates better performance.
4.4.2. Efficiency Metric
We evaluate the efficiency of models on both graph classification and regression tasks based on the training time (s), inference time (s), memory usage (MB) in training and inference phases.
4.5. Hyperparameter and Hardware Setting
For classification datasets, we set the batch size as 32 for four larger datasets (REDDIT, COLLAB, DD, and MolPCBA) and 128 for the other datasets. For regression datasets, we set the batch size as 256. To efficiently tune the hyperparameters for each model, we employed the Optuna framework (Akiba et al., 2019). For each model, we conducted $200$ trials using the Tree-structured Parzen Estimator (TPE) sampler and a MedianPruner to terminate unpromising trials early. The hyperparameter search spaces were defined as follows: the hidden dimension is selected from $\{64,128,256,512\}$, the learning rate from $\{1e-2,1e-3,1e-4,1e-5\}$, GNN layers from $\{1,2,3,4\}$, and the dropout rate from $\{0,0.1,0.2,0.3,0.4,0.5\}$. All models are trained for a maximum of 2000 epochs, with early stopping applied if no improvement is observed on the validation set within 50 epochs.
All experiments are executed on a CentOS 7 machine equipped with dual 10-core Intel® Xeon® Silver 4210 CPUs @ 2.20GHz, 8 NVIDIA GeForce RTX 2080 Ti GPUs (11GB each) and 256GB RAM.
4.6. Effectiveness Evaluation
Table 4. Evaluation on graph classification. The best and second performance are highlighted in bold and underlined, respectively. “Met.” means Metrics, “OOM” means out-of-memory, and “>3>3days” means training cannot be completed within 3 days.
Type Data Met. Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE PNA TopK GMT EPool ECS AK+ I2 VIB HGP MO RGC MVG GCA
SN I-B Acc 0.6840 0.7100 0.6860 0.7280 0.6860 0.7440 0.6980 0.7150 0.7290 0.7070 0.7250 0.7030 0.7310 0.6230 0.6970 0.7120
F1 0.6963 0.7117 0.6991 0.7391 0.7105 0.7625 0.7146 0.7158 0.7391 0.7014 0.7185 0.7078 0.7422 0.6321 0.6940 0.7147
I-M Acc 0.4620 0.4820 0.4650 0.4980 0.4670 0.5080 0.4750 0.4727 0.4967 OOM 0.4720 0.4650 0.5067 0.4130 0.4867 0.4833
F1 0.4393 0.4706 0.4512 0.4842 0.4435 0.4850 0.4563 0.4466 0.4839 OOM 0.4443 0.4376 0.4874 0.3867 0.4788 0.4726
RED Acc 0.9029 0.8965 0.9094 OOM 0.9280 0.9195 0.9260 OOM OOM OOM 0.8276 OOM 0.8625 OOM OOM OOM
F1 0.9123 0.9026 0.9132 OOM 0.9322 0.9263 0.9303 OOM OOM OOM 0.8298 OOM 0.8722 OOM OOM OOM
COL Acc 0.7644 0.7328 0.7406 OOM 0.7556 0.8164 0.7696 OOM OOM OOM 0.7528 0.6988 0.8278 OOM 0.7688 OOM
F1 0.7454 0.7099 0.7169 OOM 0.7395 0.7973 0.7520 OOM OOM OOM 0.7114 0.6702 0.8046 OOM 0.7320 OOM
BIO PRO Acc 0.7286 0.7250 0.7367 0.7372 0.7277 0.7440 0.7286 0.7061 0.7475 0.7071 0.7322 0.7304 0.7232 0.7015 0.7367 0.7322
F1 0.6596 0.6432 0.6617 0.6645 0.6712 0.7058 0.6598 0.5838 0.6978 0.6026 0.6641 0.6653 0.6442 0.6386 0.6993 0.6647
$D$ Acc 0.7310 0.7233 0.7542 0.7488 0.7156 0.7802 0.7343 OOM 0.7776 0.7335 0.7632 0.7598 0.7632 OOM OOM OOM
F1 0.6676 0.6555 0.6707 0.7245 0.6384 0.7252 0.6687 OOM 0.7267 0.6402 0.6845 0.6841 0.6844 OOM OOM OOM
ENZ Acc 0.4650 0.4833 0.5183 0.5233 0.4800 0.4950 0.4783 0.4617 0.5267 0.4650 0.4417 0.4667 0.5250 0.4917 0.5317 0.4817
F1 0.4775 0.4781 0.5098 0.5204 0.4794 0.4237 0.4737 0.4592 0.5279 0.4628 0.4337 0.4695 0.5309 0.4902 0.5314 0.4772
HIV Acc 0.9696 0.9701 0.9695 0.9703 0.9687 0.9691 0.9696 0.9689 0.9728 0.9703 0.9686 0.9689 0.9701 0.9686 0.9705 0.9701
F1 0.3192 0.3466 0.2991 0.3622 0.2221 0.2861 0.3213 0.2537 0.3682 0.3413 0.2203 0.2518 0.2502 0.2225 0.3628 0.3448
TOX Acc 0.5548 0.5553 0.5529 0.5561 0.5523 0.5412 0.5514 0.5463 0.5561 0.5580 0.5283 0.5350 0.5384 0.5112 0.5565 0.5558
F1-i 0.9120 0.9114 0.9096 0.9135 0.9083 0.9109 0.9101 0.9121 0.9138 0.9137 0.8993 0.9004 0.9048 0.8986 0.9126 0.9123
F1-a 0.3628 0.3601 0.3418 0.3839 0.2240 0.3452 0.3765 0.3858 0.3850 0.3815 0.2020 0.2125 0.2180 0.1981 0.3847 0.3695
CHE MUT Acc 0.8041 0.8570 0.8199 0.8588 0.8094 0.8254 0.8091 0.8091 0.8409 0.8062 0.7643 0.7706 0.7790 0.7067 0.8607 0.8842
F1 0.8583 0.8907 0.8644 0.8909 0.8590 0.8642 0.8506 0.8558 0.8770 0.8512 0.8255 0.8296 0.8447 0.8012 0.8912 0.9070
NCI Acc 0.8158 0.8154 0.8146 0.8183 0.8169 0.7662 0.8160 0.7866 0.8187 0.7603 0.7816 0.7825 0.7852 0.6995 0.7560 0.8122
F1 0.8160 0.8153 0.8149 0.8184 0.8171 0.7708 0.8168 0.7914 0.8190 0.7673 0.7827 0.7831 0.7855 0.6997 0.7557 0.8119
BAC Acc 0.6711 0.6645 0.6477 0.7074 0.6756 0.6513 0.6749 0.7125 0.6803 0.6078 0.6019 0.6689 0.6294 0.6089 0.7099 0.6729
F1 0.7154 0.7351 0.7099 0.7342 0.7264 0.7366 0.7436 0.7400 0.7382 0.7055 0.7001 0.6545 0.7077 0.7047 0.7366 0.7186
PCB Acc 0.5456 0.5466 0.5451 0.5470 0.5461 0.5447 >3>3days OOM 0.5523 0.5414 OOM OOM OOM OOM OOM OOM
Mi-F1 0.9850 0.9853 0.9839 0.9853 0.9849 0.9836 >3>3days OOM 0.9860 0.9850 OOM OOM OOM OOM OOM OOM
Ma-F1 0.1200 0.1500 0.1317 0.1517 0.1333 0.0167 >3>3days OOM 0.2140 0.2064 OOM OOM OOM OOM OOM OOM
Table 5. Evaluation on graph regression. The best and second performance are highlighted in bold and underlined, respectively.
Type Data Met. Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE PNA TopK GMT EPool ECS AK+ I2 VIB HGP MO RGC MVG GCA
MC 3-Cyc MAE 0.4396 0.3962 0.5119 0.4060 0.4290 0.4251 0.4238 0.0192 0.0022 0.0009 0.8776 0.4369 0.5410 0.4744 0.4231 0.3873
R2 0.6970 0.7546 0.8081 0.7493 0.7044 0.7107 0.5679 0.9996 0.9998 0.9999 0.1025 0.6850 0.5003 0.6062 0.7174 0.7595
4-Cyc MAE 0.2806 0.2536 0.5407 0.2509 0.2772 0.2720 0.2696 0.0148 0.0223 0.0062 0.6451 0.2745 0.5442 0.5398 0.2734 0.2204
R2 0.8229 0.8864 0.4011 0.8924 0.8331 0.8477 0.8406 0.9997 0.9990 0.9998 0.1424 0.8350 0.4032 0.4436 0.8606 0.8987
5-Cyc MAE 0.2782 0.1858 0.4608 0.2580 0.2399 0.2338 0.2659 0.0716 0.0340 0.0115 0.9021 0.2763 0.4532 0.4570 0.2180 0.1756
R2 0.8334 0.9364 0.4815 0.8866 0.8614 0.8936 0.8553 0.9862 0.9972 0.9996 0.0019 0.8190 0.6081 0.4939 0.9060 0.9408
6-Cyc MAE 0.3014 0.1895 0.4690 0.2841 0.2302 0.1788 0.2926 0.0862 0.0584 0.0368 0.8876 0.3035 0.4557 0.5027 0.1773 0.1784
R2 0.7926 0.9162 0.6189 0.8252 0.8803 0.9279 0.8073 0.9569 0.9956 0.9969 0.0021 0.7891 0.6219 0.6077 0.9301 0.9223
7-Cyc MAE 0.4011 0.2111 0.5898 0.2237 0.2852 0.1569 0.3942 0.1560 0.0563 0.0488 0.8273 0.4220 0.5713 0.5836 0.1500 0.2045
R2 0.5887 0.8635 0.4594 0.8469 0.7921 0.9354 0.6091 0.9460 0.9937 0.9946 0.0064 0.5487 0.4883 0.4692 0.9404 0.8770
8-Cyc MAE 0.4757 0.2632 0.5285 0.2841 0.2910 0.1377 0.4701 0.1153 0.0493 0.0399 0.7430 0.4790 0.5340 0.4808 0.1286 0.2592
R2 0.3570 0.7220 0.2212 0.7152 0.7553 0.9428 0.3840 0.9703 0.9949 0.9963 0.0330 0.3520 0.2051 0.3319 0.9466 0.7265
4-Path MAE 0.7154 0.4273 0.7340 0.2940 0.5273 0.1606 0.6361 0.0243 0.0150 0.0083 0.7784 0.5019 0.7316 0.7270 0.1512 0.4093
R2 0.1574 0.6349 0.0898 0.8619 0.5270 0.9459 0.3188 0.9921 0.9996 0.9998 0.0279 0.5390 0.0909 0.1310 0.9534 0.6579
5-Path MAE 0.6848 0.3949 0.6669 0.3628 0.5189 0.1559 0.6133 0.0134 0.0153 0.0092 0.7510 0.4846 0.6933 0.7053 0.1407 0.3805
R2 0.1050 0.6358 0.1675 0.6973 0.5015 0.9473 0.3254 0.9995 0.9995 0.9998 0.0283 0.5204 0.0823 0.0650 0.9562 0.6562
6-Path MAE 0.6164 0.3913 0.6594 0.3397 0.5629 0.1394 0.6930 0.0138 0.0133 0.0089 0.7571 0.4552 0.6627 0.6589 0.1299 0.3784
R2 0.2470 0.5958 0.1286 0.7429 0.4133 0.9539 0.1408 0.9995 0.9996 0.9998 0.0250 0.5459 0.1282 0.1280 0.9607 0.6076
4-Cliq MAE 0.3432 0.3449 0.3498 0.2504 0.3430 0.3427 0.3578 0.0090 0.0092 0.0005 0.3866 0.2355 0.3422 0.3424 0.1797 0.3401
R2 0.1612 0.1348 0.1086 0.8230 0.1213 0.1022 0.1057 0.9665 0.9964 0.9999 0.0890 0.8896 0.1604 0.1610 0.9002 0.1644
Tailed Tri MAE 0.3506 0.2886 0.3812 0.3470 0.3465 0.3665 0.3405 0.0189 0.0146 0.0034 0.8934 0.3791 0.3891 0.4095 0.3277 0.2747
R2 0.7884 0.8606 0.7303 0.7792 0.7813 0.7673 0.8053 0.9989 0.9995 0.9999 0.0176 0.7336 0.7008 0.6245 0.8177 0.8662
Chor. Cyc MAE 0.4377 0.3683 0.4390 0.3500 0.4216 0.3690 0.4246 0.0295 0.0336 0.0044 0.8720 0.4441 0.4413 0.4728 0.3671 0.3608
R2 0.6346 0.7109 0.6002 0.8121 0.6414 0.6848 0.6421 0.9970 0.9951 0.9999 0.0333 0.5746 0.5841 0.5684 0.7293 0.7149
Tri. Rec. MAE 0.4657 0.4034 0.4794 0.4410 0.4736 0.4068 0.4890 0.3619 0.3464 0.3271 0.8263 0.4286 0.4755 0.5409 0.4301 0.3943
R2 0.5211 0.6291 0.5335 0.5598 0.4921 0.6235 0.4702 0.6895 0.7063 0.7241 0.0338 0.5828 0.5037 0.3852 0.5691 0.6425
Table 6. Efficiency evaluation. We report total training (Train (s)) and inference (Infer (s)) times with two decimal places, and GPU memory usage for training (T-Mem (MB)) and inference (I-Mem (MB)) in integers. “OOM” indicates out-of-memory.
Type Data Met. Node-based Pooling-based Subgraph-based GL-based SSL-based
GCN GIN SAGE PNA TopK GMT EPool ECS AK+ I2 VIB HGP MO RGC MVG GCA
SN RED Train(s) 192.44 182.98 179.55 OOM 218.03 758.22 8365.60 OOM OOM OOM OOM OOM 371.86 OOM OOM OOM
Infer(s) 28.42 26.21 26.04 OOM 32.95 30.45 1953.91 OOM OOM OOM OOM OOM 79.35 OOM OOM OOM
T-Mem(MB) 287 288 310 OOM 287 2120 360 OOM OOM OOM OOM OOM 604 OOM OOM OOM
I-Mem(MB) 190 122 135 OOM 190 244 194 OOM OOM OOM OOM OOM 289 OOM OOM OOM
BIO ENZ Train(s) 26.05 26.34 25.23 72.33 31.95 38.85 375.25 312.46 350.25 470.05 120.48 67.50 75.29 495.23 57.53 125.45
Infer(s) 5.19 5.33 5.04 17.42 6.84 7.17 76.30 68.74 68.93 94.92 36.12 19.20 28.27 85.11 7.76 5.07
T-Mem(MB) 83 78 77 289 83 228 103 582 2369 1656 323 186 506 282 324 272
I-Mem(MB) 42 36 36 96 42 46 45 270 879 790 152 95 194 54 58 42
CHE MUT Train(s) 10.78 11.03 10.05 24.53 13.50 15.96 60.13 52.05 88.70 67.33 44.37 35.32 39.99 80.49 25.15 20.72
Infer(s) 3.86 3.89 3.51 7.94 5.27 4.01 17.74 18.91 21.18 25.10 15.80 11.63 15.18 29.78 7.64 3.94
T-Mem(MB) 42 44 44 152 42 82 46 422 775 1306 56 48 431 134 140 104
I-Mem(MB) 19 20 21 82 19 24 19 245 210 722 26 27 74 24 25 19
MC 4-Cyc Train(s) 166.50 160.95 140.88 458.70 180.57 1224.39 482.95 1180.51 645.97 1487.66 395.71 284.34 394.60 1770.94 300.21 304.95
Infer(s) 256.17 241.02 235.40 590.20 279.43 2058.62 1033.68 1776.91 920.13 2219.63 549.60 441.95 583.55 2675.22 461.87 262.40
T-Mem(MB) 88 85 85 228 88 380 104 1214 1734 3080 262 269 470 238 237 165
I-Mem(MB) 64 62 62 209 64 260 68 342 608 726 104 152 270 103 96 64
4.6.1. Graph Classification Tasks
As shown in Table 4, node-based GNNs, such as GCN and SAGE, cannot achieve satisfactory performance. These models learn node representations and use global pooling. Global pooling overlooks local structural information, which is critical for distinguishing graphs in bioinformatics (e.g., ENZYMES) and chemistry (e.g., MUTAG). PNA achieves comparatively good performance as it captures richer neighborhood distributions by leveraging multiple aggregators and degree-scalers. Secondly, pooling-based approaches, such as GMT and TopK, achieve competitive performance, particularly on social networks like COLLAB and REDDIT. These methods progressively reduce the graph size by grouping nodes or retaining only the most important ones, thereby preserving multi-level structural information. However, they are less effective on datasets where fine-grained local structures (e.g., motifs) play a critical role, such as ENZYMES and NCI1. Subgraph-based methods, such as ECS, AK+, and I2, demonstrate highly competitive performance on bioinformatics and chemistry datasets because they break graphs into meaningful substructures, enabling them to capture important patterns that other methods often miss. However, they are out-of-memory (OOM) on large graphs or high node counts, such as REDDIT and COLLAB. Fourthly, graph learning-based approaches, such as MO and HGP, perform well on noisy social datasets, such as IMDB-B and IMDB-M. By dynamically reconstructing graph structures and removing irrelevant edges or nodes, these approaches enhance robustness and improve generalization. However, they are less effective on datasets where the original graph structure is already well-formed molecular graphs, such as MUTAG. Lastly, SSL-based methods like MVG and GCA perform robustly across multiple datasets, leveraging pretraining on unlabeled graphs to learn generalizable representations. However, SSL-based approaches are computationally expensive due to graph augmentation and can lead to OOM issues.
4.6.2. Graph Regression Tasks
For graph regression tasks, the focus is primarily on evaluating how well GNNs can capture the semantics and key structural patterns of graphs, such as cycle and path counts for each graph. Lower MAE and higher $R^{2}$ scores reflect better performance. As shown in Table 5, node-based methods (excluding GIN and PNA), pooling-based models, GL-based techniques, and SSL-based approaches generally fail to deliver satisfactory results. This is because these methods are not specifically designed to enhance the expressiveness of GNNs, meaning they cannot effectively differentiate between isomorphic graphs or graphs with identical cycles. In contrast, GIN, PNA, and subgraph-based GNNs explicitly aim to improve the theoretical expressiveness of GNNs, resulting in superior performance. GIN and PNA aim to improve the theoretical expressiveness to approximate the Weisfeiler-Lehman (1-WL) isomorphism test. Subgraph-based models like ECS, AK+, and I2 consistently outperform other approaches on almost all regression datasets. By breaking graphs down into overlapping or rooted subgraphs, these methods capture rich structural details that enhance their theoretical expressivity. As a result, they are better able to distinguish between graph isomorphism classes and accurately identify important motifs, which leads to improved performance. Lastly, as the complexity of the target motif increases (e.g., from 3-Cycle to 8-Cycle), most GNNs (except subgraph-based ones) experience a drastic performance drop. This highlights the limitations of many GNN architectures in capturing higher-order dependencies and complex structural relationships.
(a) IMDB-M
(b) ENZYMES
(c) NCI1
(d) MUTAG
Figure 3. Robustness evaluation on noisy graphs.
(a) IMDB-M
(b) ENZYMES
(c) NCI1
(d) 4-Cycle
Figure 4. Imbalance data evaluation.
(a) IMDB-M
(b) ENZYMES
(c) NCI1
(d) 4-Cycle
Figure 5. Few-shot evaluation.
4.7. Efficiency Evaluation
We evaluate the efficiency of GNNs in terms of total training time, inference time, GPU peak memory usage during training and inference. As shown in Table 6, node-based GNNs exhibit the highest efficiency across all datasets. Their simplicity in aggregating neighbors’ information without additional graph processing ensures minimal time and memory usage. However, this efficiency comes at the cost of limited expressiveness as discussed in Section 4.6. Secondly, pooling-based methods strike a balance between efficiency and performance. TopK is efficient due to its node-pruning strategy, which reduces graph size while preserving key features. However, more advanced pooling approaches, such as GMT, are computationally expensive due to their clustering operations based on node similarity. Thirdly, subgraph-based GNNs (e.g., ECS, AK+, I2) are computationally intensive because they rely on generating multiple subgraphs for each graph. Despite their high resource requirements, these methods excel at capturing graph motifs and complex structural patterns, as shown in Section 4.6. Their computational cost makes them less efficient for real-time or large-scale applications. GL-based methods (e.g., VIB, HGP, and MO) dynamically reconstruct graph structures during training, which adds significant computational overhead. While they improve robustness to noise and graph quality, their iterative optimization process limits scalability to larger datasets. SSL-based methods (e.g., RGC, MVG, GCA) are computationally expensive during training because they require graph augmentations and additional contrastive learning stage.
In summary, node-based GNNs are the most efficient but lack expressiveness, while pooling-based models strike a balance between efficiency and performance. Subgraph-based and GL-based approaches offer superior expressiveness but suffer from high computational costs. SSL-based methods, though computationally expensive during training, are efficient during inference, making them suitable for pretraining scenarios.
4.8. More Scenarios
We further evaluate GNNs in three realistic and challenging scenarios, including noisy graphs, imbalanced graphs, and few-shot graphs. To ensure clarity, we select one representative graph dataset from each domain for evaluation, i.e., IMDB-M, ENZYMES, NCI1, and 4-Cycle. Additionally, we compare some of the most representative models from each GNN category, including GCN (node-based), GMT (pooling-based), and GCA (SSL-based). For top perform, we select AK+ (subgraph-based), MO (GL-based) for classification tasks, while I2 (subgraph-based) and HGP (GL-based) for regression tasks.
4.8.1. Robustness Evaluation on Noisy Graphs
[…]
4.8.2. Imbalance Data Evaluation
For the graph classification task, given a dataset $\mathcal{G}={(G_{i},y_{i})}$, we simulate class imbalance by adjusting the proportion of training samples in each class to follow the sequence $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}},\dots,\frac{1}{|\mathcal{Y}|^{\beta}}\}$, where $\beta\in\{0,0.5,1,1.5,2\}$ determines the degree of imbalance. The sample count for the first class remains constant across all $\beta$ settings. For graph regression tasks, where $y_{i}$ spans the interval $y_{\text{min}},y_{\text{max}}]$, we divide the label range into three equal-sized buckets and set the group proportions as $\{1,\frac{1}{2^{\beta}},\frac{1}{3^{\beta}}\}$ to create imbalance, with larger $\beta$ corresponding to higher level of imbalance. As shown in Figure 4, accuracy drops across various GNN architectures as imbalance increases, particularly in datasets with minority classes. Node-based GNN (e.g., GCN) and pooling-based GNN (GMT) struggle because their global aggregation mechanisms tend to average out minority signals, making it difficult to distinguish rare patterns. Subgraph-based (AK+) and graph learning-based (MO) GNNs, while capturing richer structural information, do not explicitly address class imbalance, leading to performance degradation under extreme imbalance, especially for less classes. SSL-based model (GCA), although capable of leveraging abundant unlabeled data, likewise fails to correct imbalance on its own and converges to performance similar to other baselines unless augmented with imbalance-aware strategies (Liu et al., 2023).
4.8.3. Few-shot Evaluation
We simulate data scarcity by limiting the number of training samples. For classification tasks, we construct training sets where the number of labeled graphs per class is in $\gamma\in\{5,10,15,20,25\}$. For regression tasks, we partition the label into five equal-width buckets and uniformly sample $\gamma$ training instances per bucket. The results, as summarized in Figure 5, reveal that most GNN architectures exhibit significant performance degradation as $\gamma$ decreases. Node-based (GCN) and pooling-based (GMT) models are particularly affected, as their global aggregation mechanisms rely heavily on abundant labeled data to learn meaningful representations. Interestingly, while subgraph-based models such as AK+ and I2, as well as graph learning-based models like MO and HGP, are theoretically capable of leveraging local structural patterns, they do not demonstrate substantial resilience to data scarcity in practice. This is primarily because current implementations lack explicit mechanisms to identify and prioritize the most informative subgraphs or adaptively focus on critical features when labeled data is limited. Their performance improvements over standard baselines are thus marginal in few-shot scenarios, indicating that richer local modeling alone does not guarantee data efficiency (Wang et al., 2023c).
4.9. Correlation Analysis
4.9.1. Graph Topological Features
To investigate the intrinsic relationship between graph topology and GNN model performance, we characterize each graph using a comprehensive set of topological features. These features capture properties ranging from macroscopic global characteristics to fine-grained local patterns. We categorize them into three groups: (i) Global Structural Features, (ii) Local Structural Features, and (iii) Node Distributions.
Global Structural Features. These metrics provide a high-level summary of the entire graph’s connectivity and organization. They typically return a single scalar value for each graph.
  • Average Degree (Deg.) and Density (Den.). For a graph $G=(V,E)$, let $|V|$ and $|E|$ be the number of nodes and edges, respectively. $\text{Avg. Deg}=\frac{2|E|}{|V|}$. Density is the ratio of actual edges to possible edges:
    (16) $\text{Density}=\frac{2|E|}{|V|(|V|-1)}.$
    They measure the overall sparsity.
  • Average Shortest Path Length (SPL) and Diameter (Dia.). Let $d(u,v)$ denote the shortest path distance between nodes $u$ and $v$. $\text{SPL}=\frac{1}{|V|(|V|-1)}\sum_{u\neq v}d(u,v)$, and $\text{Diameter}=\max_{u\neq v}d(u,v)$. For disconnected graphs, these are computed on the largest connected component.
  • Degree Assortativity (Assr.). Calculated as the Pearson correlation coefficient between the degrees of connected nodes. Positive values indicate that high-degree nodes tend to connect with other high-degree nodes.
  • Modularity (Mod.). Quantifies the strength of community structure. A positive value indicates that connections within communities are denser than expected by random chance.
Local Structural Features. These metrics focus on small, recurring substructures. For node-level metrics, we aggregate them into graph-level features.
  • Average Clustering Coefficient (CC). For a node $v$, the local clustering coefficient $C_{v}$ is the density of the subgraph induced by its neighbors. We report the average over all nodes: $\text{Avg. CC}=\frac{1}{|V|}\sum_{v\in V}C_{v}$.
  • Triangle Counts (Tri.). Let $T_{v}$ be the number of triangles node $v$ participates in. Unlike global metrics, to capture both the prevalence and distribution of triangles, we report both the mean ($\mu$) and standard deviation ($\sigma$) of $T_{v}$ across all nodes. High $\mu$ implies a dense graph, while high $\sigma$ indicates a significant degree of heterogeneity.
Node Distributions. These features analyze the distribution of node importance. Since these are originally defined at the node level, we characterize the entire graph by reporting both the Mean ($\mu$) and Standard Deviation ($\sigma$) for each metric.
  • Betweenness Centrality (BC.). $C_{B}(v)$ represents the fraction of all-pairs shortest paths passing through node $v$. We report $\mu_{\text{BC}}$ and $\sigma_{\text{BC}}$. A high $\sigma_{\text{BC}}$ indicates the presence of critical ”bridge” nodes controlling information flow.
  • PageRank (PR.). Measures node influence based on transitive importance. We report $\mu_{\text{PR}}$ and $\sigma_{\text{PR}}$. A high $\sigma_{\text{PR}}$ reveals a hierarchical structure with distinct authority nodes.
  • K-Core Number (KC.). The core number $k_{v}$ is the largest $k$ such that $v$ belongs to a $k$-core. We report $\mu_{\text{KC}}$ and $\sigma_{\text{KC}}$. A high $\mu_{\text{KC}}$ indicates global robustness, while high $\sigma_{\text{KC}}$ reflects a core-periphery structure.
4.9.2. Evaluation
Given GNNs $\mathcal{M}=\{M_{i}\}_{i=1}^{n}$ and graph dataset $\mathcal{D}=\{G_{j}\}_{j=1}^{m}$, we construct a performance matrix $\mathbf{P}\in\mathbb{R}^{n\times m}$ where $\mathbf{P}[i][j]$ is the result of $M_{i}$ on $G_{j}$ (e.g., accuracy or MAE). We denote each $k$-th topological feature of $\mathcal{D}$ as $\mathbf{f}_{k}$, where $\mathbf{f}_{k}[j]$ is the value of the $k$-th feature for graph $G_{j}$. We assess the relationship between model performance and $k$-th feature with Spearman’s correlation $\rho(\mathbf{P}[i],\mathbf{f}_{k})$, with a significance threshold of $\alpha=0.05$.
4.9.3. Insights
Figure 6. Correlation between graph features and GNNs. Each cell shows the Spearman’s $\rho$ with * for $p<0.05$ and $\dagger$ for $0.05\leq p<0.1$.
As illustrated in Figure 6, we observe:
  • Structurally simpler models (e.g., Node-based) show weak correlations with most features, reflecting their reliance on local aggregation rather than global topology.
  • Graph density negatively correlates with most models, suggesting high connectivity exacerbates over-smoothing and noise.
  • High local clustering and assortativity negatively impact Subgraph-based and SSL-based models, likely due to the “rich-club” effect distracting from peripheral structures.
  • Conversely, graph sparsity and high Betweenness Centrality positively correlate with Hierarchical (e.g., HGP) and SSL-based models, demonstrating their ability to leverage clear structures for long-range information flow.
Most importantly, the absence of universally significant correlations confirms that (i) model selection cannot rely on a single structural indicator, and (ii) no single architecture consistently dominates. This underscores the critical importance of comprehensive benchmarks across diverse domains, as conducted in OpenGLT, to provide an empirical basis for scenario-specific model selection.
5. Conclusion and Future Direction
This paper presents a comprehensive experimental study on graph neural networks for graph-level tasks, categorizing existing models, introducing a unified evaluation framework (OpenGLT), and benchmarking 16 GNNs across 26 diverse datasets under challenging scenarios such as noise, imbalance, and limited data. Our results reveal that no single model excels universally, highlighting important trade-offs between expressiveness and efficiency, and emphasizing the need for robust evaluation in realistic conditions.
Based on these findings, promising future directions include the development of scenario-adaptive or hybrid GNN architectures that dynamically leverage different model strengths, research into lightweight and scalable algorithms for practical deployment, and the incorporation of transfer and foundation model techniques to improve generalization and data efficiency, particularly when labeled data is scarce. Our study provides valuable benchmarks and guidance for advancing GNN research on graph-level tasks.